Trapezoid ABCD has parallel sides AB and CD, a right angle at D, and the lengths AB = 15. BC = 10. and CD = 7. find the lenghth DA.

thankks!

Answer by orca(409)

Draw CE perpendicular to AB.
In right triangle EBC, BC = 10, EB = 15 - 7 = 8, so EC = sqrt(10^2 - 8^2) = 6
As quadrilateral ADCE is a rectangle, so DA = CE.
Thus DA = EC = 6